Van Rushing Hunting Goods’ fiscal year ends on December 31. At the end of the 2021 fiscal year, the company had notes payable of $17.8 million due on February 8, 2022. Rushing sold 5.0 million shares of its $0.25 par, common stock on February 3, 2022, for $15.0 million. The proceeds from that sale along with $2.8 million from the maturation of some 3-month CDs were used to pay the notes payable on February 8. Through his attorney, one of Rushing’s construction workers notified management on January 5, 2022, that he planned to sue the company for $1 million related to a work-site injury on December 20, 2021. As of December 31, 2021, management had been unaware of the injury, but reached an agreement on February 23, 2022, to settle the matter by paying the employee’s medical bills of $78,000. Rushing’s financial statements were finalized on March 3, 2022. Required: 1. What amount(s) if any, related to the situations described should Rushing report among current liabilities in its balance sheet at December 31, 2021? 2. What amount(s) if any, related to the situations described should Rushing report among long-term liabilities in its balance sheet at December 31, 2021? 3. What amount(s) if any, related to the situations described should Rushing report among current liabilities and long-term liabilities in its balance sheet at December 31, 2021 if the settlement agreement had occurred on March 15, 2022, instead? 4. What amount(s) if any, related to the situations described should Rushing report among current liabilities and long-term liabilities in its balance sheet at December 31, 2021 if the work-site injury had occurred on January 3, 2022, instead?
Answer #1

Requirement 1:

Current liabilities = portion of notes payable due in 3 months + medical bills

Current liabilities = [ 17,800,000 - 15,000,000] + 78000 = 2,800,000 +78000 = 2,878,000

Requirement 2:

Long term liabilities = portion of notes payable due as through stock

Long term liabilities = 17,800,000 - 2,800,000 = 15,000,000

Requirement 3:

If settlement occurs on March 15, then medical bills will not be included and amount will be as follows. This is because the financial statements were finalized on 3rd march.

Current liabilities = [ 17,800,000 - 15,000,000] = 2,800,000

Long term liabilities = 15,000,000

Requirement 4:

If work site injury occurs on January 3 2022, then medical bills will not be included and amount will be as follows. This is because the injury happened next year and also the sue was not confirmed and probable in nature

Current liabilities = [ 17,800,000 - 15,000,000] = 2,800,000

Long term liabilities = 15,000,000